OpenStudy (albert0898):

HELP! Explain the steps to me please. I know the first step is to distribute the equation on the right. If 6x^2 + x - c = (cx - 1)(3x + 2), what is the value of c? F. -2 G. -1 H. 2 J. 3 K. 6

OpenStudy (albert0898):

\[6x^2 + x - c = (cx - 1)(3x + 2)\]

OpenStudy (astrophysics):

So go ahead and distribute

OpenStudy (albert0898):

\[6x^2 + x - c = 3cx^2 + 2cx - 3x - 2\]

OpenStudy (zale101):

Combine like terms.

OpenStudy (albert0898):

\[6x^2 + 4x - c = 2cx - 2\]

OpenStudy (astrophysics):

You can factor out c, move the -c on the right I guess

OpenStudy (albert0898):

So... \[6x^2 + 4x = 2cx + c - 2\] What's next?

OpenStudy (astrophysics):

move the -2 to the left

OpenStudy (astrophysics):

factor out c

OpenStudy (astrophysics):

solve for c

OpenStudy (albert0898):

\[6x^2 + 4x + 2 = 2cx + c\] \[6x^2 + 4x + 2 = c(2x + 1)\]

OpenStudy (zale101):

(Divide both sides by (2x+1)

OpenStudy (albert0898):

Oh boy... I just learned this in precalc. I hate this topic although I got a 95.

imqwerty (imqwerty):

d00d no need to distribute u know that the 6x^2 term will be made when 3x will get multiplied to cx so \[3x \times cx=6x^2\] x=2 (;

OpenStudy (astrophysics):

\(6x^2+x-c=(cx-1)(3x+2)\) \(6x^2+x-c = 3cx^2+2cx-3x-2\) \(6x^2+x-c - 3cx^2-2cx+3x+2=0\) \(6x^2+4x+2-3cx^2-2cx-c=0\) \(6x^2+4x+2=3cx^2+2cx+c\) \(6x^2+4x+2=c(3x^2+2x+1)\) \[c = \frac{ 6x^2+4x+2 }{ 3x^2+2x+1 }\] sorry I'm slow with latex

OpenStudy (astrophysics):

Changing all the square brackets to round

OpenStudy (astrophysics):

I sort of overcomplicated this I guess, but yeah x = 2 -.- qwerty

OpenStudy (albert0898):

Astro, I understand it a little better now. Qwerty, can you explain your method a little clearer?

OpenStudy (astrophysics):

All methods work

imqwerty (imqwerty):

\(\Huge\ddot\smile\) mmk

OpenStudy (astrophysics):

\[c = \frac{ 6x^2+4x+2 }{ 3x^2+2x+1 } \implies \frac{ 2(3x^2+2x+1) }{ (3x^2+2x+1) } = 2\]

imqwerty (imqwerty):

so we have this- \(6x^2+x-c=(cx-1)(3x+2)\) We gotta prove that the left hand side is equal to the right hand side now we got only 1term of the degree 2 on the Left and it is \(\huge\color{red}{6x^2}\) and because both sides gotta be equal we need a 6x^2 on the right as well now look at this-> right hand side->\(\large\color{violet}{(cx-1)(3x+2)}\) now when we open the brackets up we need to get a \(\color{red}{6x^2}\) which is a 2nd degree term a second degree term is obtained when we multiply a 1st degree term with another 1st degree term now look at this->\((\color{blue}{cs}-1)(\color{blue}{3x}+2)\) here \(\color{gold}{cs~and~3x}\) are the only 1st degree terms multiplying them together we must get a \(\color{red}{6x^2}\) so that LHS=RHS so \(\large\color{black}{cs \times 3x=6x^2}\) \(\large\color{black}{c=2}\)

OpenStudy (albert0898):

I understand it now. Thank you to all!

imqwerty (imqwerty):

ayy np (;